OpenStudy (anonymous):

What mass of mercury would require 65.5 J of energy to raise the temerature of 20°C to 100°C? The specific heat of mercury is 0.140 J/g*°C.

OpenStudy (anonymous):

Answer: 5.8 g 8.7 g 730 g 0.17 g

OpenStudy (anonymous):

@aaronq

OpenStudy (mathstudent55):

Q = mc(dT) 65.5 ~J = m (0.140 J/g*°C)(80 °C) \(m = \dfrac{65.5 J}{(0.140 \frac{J}{g~°C})(80 °C)}\)